OpenStudy (anonymous):

The volume and surface area of a cube are calculated [Keep in mind, this means that length=width=height]. The length is then doubled, while the width and height are kept the same. This results in an increased volume by 64 cubic centimeters. The surface area increased from 96 square centimeters to 160 square centimeters. What where the original dimensions?

OpenStudy (king):

4?

OpenStudy (anonymous):

How did you get that?

OpenStudy (ash2326):

earlier l,b,h are a,a,a repetively Volume1=a^3 now length =2a Volume2=2a^3 Volume2=64+Volume1 2a^3=64+a^3 a^3=64 so a =4 so l=b=h=4